# Can the diffeomorphism and Gauss constraints be holonomy corrected in the deformed algebra approach to modified gravity?

### v1

@@ III.3.1 $\tilde{g}=g$
To fulfill this condition, either $g=\mathfrak{c}+m_{1}(\mathfrak{p})$, with $m_{1}$ an arbitrary function of $\mathfrak{p}$ vanishing at the classical limit, or $\xi_{3}$ has to fulfill Eq. (41). It should be underlined that the usual correction does not match this requirement, which makes the situation rather “unnatural”. Of course, $g=\tilde{g}$ must satisfy Eq. (42). The other corrections ($h$ and $l$) remain free.
@@ III.3.3 Case $\xi_{3}=\chi_{2}=0$
This subsection focuses on the case where the two counterterms $\xi_{3}$ and $\chi_{2}$ are both vanishing. This allows to underline the connection between the different corrections.
@@
The corrections appearing in the Gauss and diffeomorphism constraints must be proportional one to the other. In particular, in the case where the Gauss constraint is not corrected, Eq. (51) leads to $h=\mathfrak{c}/L$. In order to get the classical limit of $h$, $L$ must be equal to one, leading to $h=\mathfrak{c}$. Thus, there cannot be a diffeomorphism correction without correcting the Gauss constraint (the reciprocal is also true). Furthermore, in this case, $L=1$ implies that $\chi_{1}=0$ through Eq. 38.
@@
If $\tilde{g}\neq g\partial_{\mathfrak{c}}g$, this implies $h=\mathfrak{c}+m_{2}(\mathfrak{p})$, where $m_{2}$ is an arbitrary function of $\mathfrak{p}$ which vanishes at the classical limit ($\mathfrak{p}\rightarrow+\infty$). This leads to Eq. (51) implying that $l=\mathfrak{c}+m_{2}(\mathfrak{p})$, as $L=1$ for the same reason as earlier (leading to $\chi_{1}=0$).
However, if $\tilde{g}=g\partial_{\mathfrak{c}}g$, then the diffeomorphism and Gauss constraints are free, as long as they are proportional to each other, see Eq. (51). This case is particularly noteworthy since, if $\tilde{g}=g$, then Eq. (40) yields $g=\mathfrak{c}+m_{3}(\mathfrak{p})$, with $m_{3}$ a function of $\mathfrak{p}$ which vanishes at the classical limit.
@@ IV Summary and conclusion
If only the diffeomorphism and Hamiltonian constraints are corrected, the cancellation of anomalies implies that:
@@
If all the constraints are corrected, then the cancellation of the anomalies implies that :
@@
  • equation (53) is not enough to get a first-class algebra. It is the condition (42) that needs to be fufiled by $g$.
@@
However, if the Gauss constraint is also modified, a first class Poisson algebra can be obtained. Interestingly, the differential equations for the correction functions are not the usual ones.
@@
Appendix A List of anomalies without correcting the Gauss constraint
@@
$\mathcal{A}_{1}^{\left\{\mathbb{D},\mathbb{G}\right\}}=(1+\xi_{2})\mathfrak{c}-h-\xi_{3}.$ (A.54)
For the $\left\{\mathbb{H},\mathbb{G}\right\}$ bracket :
$\mathcal{A}_{1}^{\left\{\mathbb{H},\mathbb{G}\right\}}=g^{2}-2\mathfrak{c}(\tilde{g}+\alpha_{6})+(\tilde{g}^{2}+\alpha_{7})+4\mathfrak{p}g\partial_{\mathfrak{p}}g$ (A.55)
$\mathcal{A}_{2}^{\left\{\mathbb{H},\mathbb{G}\right\}}=\mathfrak{c}(1+\alpha_{4})+(\tilde{g}+\alpha_{6})-2g\partial_{\mathfrak{c}}g$ (A.56)
$\mathcal{A}_{3}^{\left\{\mathbb{H},\mathbb{G}\right\}}=-\beta_{10}$ (A.57)
$\mathcal{A}_{4}^{\left\{\mathbb{H},\mathbb{G}\right\}}=\beta_{12}.$ (A.58)
@@
$\mathcal{A}_{1}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(\tilde{g}^{2}+\alpha_{2})(\xi_{2}-2-3\xi_{1})-4(\tilde{g}+\alpha_{1})(h+\xi_{3})+6g^{2}$ (A.59)
$\mathcal{A}_{2}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(1+\alpha_{3})(\xi_{1}-\xi_{2})$ (A.60)
$\mathcal{A}_{3}^{\left\{\mathbb{H},\mathbb{D}\right\}}=2(1+\alpha_{4})(h+\xi_{3})+2(\tilde{g}+\alpha_{6})(1+\xi_{2})-4g(1+\xi_{2}+\mathfrak{p}\partial_{\mathfrak{p}}\xi_{2})\partial_{\mathfrak{c}}g+(g^{2}+4\mathfrak{p}g\partial_{\mathfrak{p}}g)\partial_{\mathfrak{c}}\xi_{2}$ (A.61)
$\mathcal{A}_{4}^{\left\{\mathbb{H},\mathbb{D}\right\}}=4g(1+\xi_{1}+\mathfrak{p}\partial_{\mathfrak{p}}\xi_{1})\partial_{\mathfrak{c}}g-(g^{2}+4\mathfrak{p}g\partial_{\mathfrak{p}}g)\partial_{\mathfrak{c}}\xi_{1}-2(1+\alpha_{5})(h+\xi_{3})-2(\tilde{g}+\alpha_{6})(1+\xi_{1})$ (A.62)
$\mathcal{A}_{5}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(\tilde{g}^{2}+\alpha_{7})(1+\xi_{2})-2(\tilde{g}+\alpha_{6})(h+\xi_{3})-4\mathfrak{p}g(\partial_{\mathfrak{c}}g)\partial_{\mathfrak{p}}(h+\xi_{3})+(g^{2}+4\mathfrak{p}g\partial_{\mathfrak{p}}g)\partial_{\mathfrak{c}}(h+\xi_{3})$ (A.63)
$\mathcal{A}_{6}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(\tilde{g}^{2}+\alpha_{8})(2+3\xi_{1}-\xi_{2})-2(\tilde{g}^{2}+\alpha_{7})(1+\xi_{1})$ (A.64)
$\mathcal{A}_{7}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(1+\alpha_{9})(\xi_{1}-\xi_{2}).$ (A.65)
$\mathcal{A}_{8}^{\left\{\mathbb{H},\mathbb{D}\right\}}=\partial_{\mathfrak{c}}\xi_{1}$ (A.66)
$\mathcal{A}_{9}^{\left\{\mathbb{H},\mathbb{D}\right\}}=\partial_{\mathfrak{c}}\xi_{2}$ (A.67)
$\mathcal{A}_{10}^{\left\{\mathbb{H},\mathbb{D}\right\}}=1+2\beta_{1}-(1+\xi_{1})(1+\beta_{2})$ (A.68)
$\mathcal{A}_{11}^{\left\{\mathbb{H},\mathbb{D}\right\}}=1+\beta_{5}-(1+\xi_{1})(1+\beta_{6})$ (A.69)
$\mathcal{A}_{12}^{\left\{\mathbb{H},\mathbb{D}\right\}}=\partial_{\mathfrak{c}}(h+\xi_{3})-(1+\xi_{2})(1+\beta_{10})$ (A.70)
$\mathcal{A}_{13}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(1+\xi_{2})(1+\beta_{12})-\partial_{\mathfrak{c}}(h+\xi_{3})$ (A.71)
$\mathcal{A}_{14}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(1+\xi_{1})(2+\beta_{10}+\beta_{9})-2-2\beta_{6}$ (A.72)
$\mathcal{A}_{15}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(1+\xi_{1})(\beta_{11}-\beta_{12})$ (A.73)
$\mathcal{A}_{16}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(1+\xi_{1})(1+\beta_{8})-1$ (A.74)
$\mathcal{A}_{17}^{\left\{\mathbb{H},\mathbb{D}\right\}}=(1+\xi_{1})(1+\beta_{4})-1.$ (A.75)
Finally the $\left\{\mathbb{H},\mathbb{H}\right\}$ bracket :
$+(\tilde{g}^{2}+\alpha_{2})(2+3\alpha_{5}-\alpha_{4})$ (A.76)
$+2(\tilde{g}+\alpha_{1})(\tilde{g}^{2}-2\alpha_{7}+3\alpha_{8})$ (A.77)
$-(g^{2}+4\mathfrak{p}g\partial_{\mathfrak{p}}g)\partial_{\mathfrak{c}}\alpha_{3}$ (A.78)
$\mathcal{A}_{4}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $(1+\alpha_{3})\left[(1+\xi_{2})(1+\alpha_{5})-1-\alpha_{4}\right]$ (A.79)
$\mathcal{A}_{5}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $-\xi_{1}(1+\alpha_{3})(1+\alpha_{5})$ (A.80)
$\mathcal{A}_{6}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\partial_{\mathfrak{c}}\beta_{1}$ (A.81)
$\mathcal{A}_{7}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\partial_{\mathfrak{c}}\beta_{3}$ (A.82)
$\mathcal{A}_{8}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $-\partial_{\mathfrak{c}}\beta_{2}$ (A.83)
$\mathcal{A}_{9}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $-\partial_{\mathfrak{c}}\beta_{4}$ (A.84)
$\mathcal{A}_{10}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\partial_{\mathfrak{c}}(\beta_{2}+\beta_{4})$ (A.85)
$\mathcal{A}_{11}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\beta_{1}-\beta_{3}-\beta_{5}-\beta_{3}\beta_{5}$ (A.86)
$\mathcal{A}_{12}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\beta_{1}+\beta_{7}+\beta_{1}\beta_{7}-\beta_{3}$ (A.87)
$\mathcal{A}_{13}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\beta_{1}+\beta_{3}+\beta_{6}+\beta_{3}\beta_{6}+\beta_{8}+\beta_{1}\beta_{8}-\beta_{2}-\beta_{4}$ (A.88)
$\mathcal{A}_{14}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\left[1+\beta_{2}\right](2-\alpha_{4}+3\alpha_{5})-2\partial_{\mathfrak{c}}(\tilde{g}+\alpha_{1})$ (A.89)
$\mathcal{A}_{15}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\left[1+\beta_{4}\right](-2+\alpha_{4}-3\alpha_{5})+2\partial_{\mathfrak{c}}(\tilde{g}+\alpha_{1})$ (A.90)
$\mathcal{A}_{16}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $\partial_{\mathfrak{c}}\alpha_{3}$ (A.91)
$\mathcal{A}_{17}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $-6\left[\tilde{g}+\alpha_{1}\right]\left[1+\beta_{6}\right]+2g(\partial_{\mathfrak{c}}g)(3+3\beta_{1}-2\mathfrak{p}\partial_{\mathfrak{p}}\beta_{1})+g^{2}\partial_{\mathfrak{c}}\beta_{1}+4g\mathfrak{p}(\partial_{\mathfrak{p}}g)(\partial_{\mathfrak{c}}\beta_{1})$ (A.92)
$-\partial_{\mathfrak{c}}\alpha_{2}-g^{2}\partial_{\mathfrak{c}}\beta_{2}-4g\mathfrak{p}(\partial_{\mathfrak{p}}g)(\partial_{\mathfrak{c}}\beta_{2})$ (A.93)
$+\partial_{\mathfrak{c}}\alpha_{2}+g^{2}\partial_{\mathfrak{c}}\beta_{4}+4g\mathfrak{p}(\partial_{\mathfrak{p}}g)(\partial_{\mathfrak{c}}\beta_{4})$ (A.94)
$\mathcal{A}_{20}^{\left\{\mathbb{H},\mathbb{H}\right\}}=$ $3\mathfrak{p}(\tilde{g}+\alpha_{1})+3\mathfrak{p}(\tilde{g}+\alpha_{1})\beta_{8}-g(\partial_{\mathfrak{c}}g)\left[3\mathfrak{p}(1+\beta_{3})+2\mathfrak{p}^{2}\partial_{\mathfrak{p}}\beta_{3}\right]+\frac{\mathfrak{p}}{2}g^{2}\partial_{\mathfrak{c}}\beta_{3}+2\mathfrak{p}^{2}g(\partial_{\mathfrak{p}}g)(\partial_{\mathfrak{c}}\beta_{3}).$ (A.95)
Appendix B List of anomalies when correcting every constraints
$\mathcal{A}_{1}^{\left\{\mathbb{D},\mathbb{G}\right\}}=$ $\mathfrak{p}(1+\xi_{2})(l+\chi_{2})-(h+\xi_{3})(\mathfrak{p}+\chi_{1}).$ (B.1)
For the $\left\{\mathbb{H},\mathbb{G}\right\}$ bracket, the anomalies read :
$\mathcal{A}^{\{\mathbb{H},\mathbb{G}\}}_{1}$= $g^{2}\partial_{\mathfrak{c}}\chi_{1}+4\mathfrak{p}g\left[(\partial_{\mathfrak{p}}g)(\partial_{\mathfrak{c}}\chi_{1})-(\partial_{\mathfrak{c}}g)(1+\partial_{\mathfrak{p}}\chi_{1})\right]+2\mathfrak{p}(1+\alpha_{4})(l+\chi_{2})+2(\tilde{g}+\alpha_{6})(\mathfrak{p}+\chi_{1})$ (B.2)
$\mathcal{A}^{\{\mathbb{H},\mathbb{G}\}}_{2}$= $(\tilde{g}^{2}+\alpha_{7})(\mathfrak{p}+\chi_{1})-2\mathfrak{p}(\tilde{g}+\alpha_{6})(l+\chi_{2})+\mathfrak{p}g^{2}\partial_{\mathfrak{c}}(l+\chi_{2})+4\mathfrak{p}^{2}g\left[(\partial_{\mathfrak{p}}g)\partial_{\mathfrak{c}}(l+\chi_{2})-(\partial_{\mathfrak{c}}g)\partial_{\mathfrak{p}}(l+\chi_{2})\right]$ (B.3)
$\mathcal{A}^{\{\mathbb{H},\mathbb{G}\}}_{3}$= $\partial_{\mathfrak{c}}\chi_{1}$ (B.4)
$\mathcal{A}^{\{\mathbb{H},\mathbb{G}\}}_{4}$= $\mathfrak{p}\partial_{\mathfrak{c}}(l+\chi_{2})-(\mathfrak{p}+\chi_{1})(1+\beta_{10})$ (B.5)
$\mathcal{A}^{\{\mathbb{H},\mathbb{G}\}}_{5}$= $(\mathfrak{p}+\chi_{1})(1+\beta_{12})-\mathfrak{p}\partial_{\mathfrak{c}}(l+\chi_{2}).$ (B.6)
For the $\left\{\mathbb{H},\mathbb{D}\right\}$ bracket,
@@ Appendix B List of anomalies when correcting every constraints
Finally for the $\left\{\mathbb{H},\mathbb{H}\right\}$ bracket,
@@
Appendix C Summary of the counter-terms in the case of all the constraints being corrected
Here $\xi_{3}(\mathfrak{c},\mathfrak{p})$ is left arbitrary for generality purposes. First, the matter counter-terms :
$\beta_{1}=\beta_{2}=\beta_{3}=\beta_{4}=\beta_{5}=\beta_{6}=\beta_{7}=\beta_{8}=0$ (C.1)
$\beta_{9}=1-\partial_{\mathfrak{c}}(h+\xi_{3})$ (C.2)
$\beta_{10}=\beta_{11}=\beta_{12}=\partial_{\mathfrak{c}}(h+\xi_{3})-1$ (C.3)
$\beta_{13}=\alpha_{3}[K,\alpha_{9}(\mathfrak{p})]\partial_{\mathfrak{c}}(g\partial_{\mathfrak{c}}g)+\partial_{\mathfrak{c}}(g\partial_{\mathfrak{c}}g)-1.$ (C.4)
For the Hamiltonian counter-terms :
$\alpha_{1}=g\partial_{\mathfrak{c}}g-\tilde{g}$ (C.5)
$\alpha_{2}=3g^{2}-\tilde{g}^{2}-2g(h+\xi_{3})\partial_{\mathfrak{c}}g$ (C.6)
$\alpha_{3}=\frac{1}{\sqrt{\mathfrak{p}}}\left[K+\int^{p}_{1}\frac{\alpha_{9}(u)}{2\sqrt{u}}du\right]$ (C.7)
$\alpha_{4}=\alpha_{5}=(\partial_{\mathfrak{c}}g)^{2}+g\partial^{2}_{\mathfrak{c}}g-1$ (C.8)

### v2

@@ III.3.1 $\tilde{g}=g$
To fulfill this condition, either $g=\mathfrak{c}+m_{1}(\mathfrak{p})$, with $m_{1}$ an arbitrary function of $\mathfrak{p}$ vanishing at the classical limit, or $\xi_{3}$ has to fulfill Eq. (41). It should be emphasized that the usual correction does not match this requirement, which makes the situation rather “unnatural”. Of course, $g=\tilde{g}$ must satisfy Eq. (42). The other corrections ($h$ and $l$) remain free.
@@ III.3.3 Case $\xi_{3}=\chi_{2}=0$
This subsection focuses on the case where the two counterterms $\xi_{3}$ and $\chi_{2}$ are both vanishing. This allows one to emphasize the connection between the different corrections.
@@
The corrections appearing in the Gauss and diffeomorphism constraints must be proportional one to the other. In particular, in the case where the Gauss constraint is not corrected, Eq. (50) leads to $h=\mathfrak{c}/L$. To get the classical limit of $h$, $L$ must be equal to one, leading to $h=\mathfrak{c}$. Thus, there cannot be a diffeomorphism correction without correcting the Gauss constraint (the reciprocal is also true). Furthermore, in this case, $L=1$ implies that $\chi_{1}=0$ through Eq. 38.
@@
If $\tilde{g}\neq g\partial_{\mathfrak{c}}g$, this implies $h=\mathfrak{c}+m_{2}(\mathfrak{p})$, where $m_{2}$ is an arbitrary function of $\mathfrak{p}$, which vanishes at the classical limit ($\mathfrak{p}\rightarrow+\infty$). This leads to Eq. (50) implying that $l=\mathfrak{c}+m_{2}(\mathfrak{p})$, as $L=1$ for the same reason as earlier (leading to $\chi_{1}=0$).
However, if $\tilde{g}=g\partial_{\mathfrak{c}}g$, then the diffeomorphism and Gauss constraints are free, as long as they are proportional to each other; see Eq. (50). This case is particularly noteworthy since, if $\tilde{g}=g$, then Eq. (40) yields $g=\mathfrak{c}+m_{3}(\mathfrak{p})$, with $m_{3}$ a function of $\mathfrak{p}$ that vanishes at the classical limit.
@@ IV Summary and conclusion
If only the diffeomorphism and Hamiltonian constraints are corrected, the cancellation of anomalies implies the following:
@@
If all the constraints are corrected, then the cancellation of the anomalies implies the following:
@@
  • Equation (52) is not enough to get a first-class algebra. It is the condition (42) that needs to be fufilled by $g$.
@@
However, if the Gauss constraint is also modified, a first-class Poisson algebra can be obtained. Interestingly, the differential equations for the correction functions are not the usual ones.
@@
Appendix A Expression of constraints when all corrected
First, the Hamiltonian constraint reads
$\mathbb{H}\left[N\right]$ $=\mathbb{H}_{\mathfrak{g}}\left[N\right]+\mathbb{H}_{\mathfrak{m}}\left[N\right]$
$=\frac{1}{2\kappa}\int d{\bf{x}}\;\big({\bf{N}}\big[\mathcal{H}_{\mathfrak{g}}^{(0)}+\mathcal{H}_{\mathfrak{g}}^{(2)}\big]+\delta N\mathcal{H}_{\mathfrak{g}}^{(1)}\big)$
$+\frac{1}{2\kappa}\int d{\bf{x}}\;\big({\bf{N}}\big[\mathcal{H}_{\mathfrak{m}}^{(0)}+\mathcal{H}_{\mathfrak{m}}^{(2)}\big]+\delta N\mathcal{H}_{\mathfrak{m}}^{(1)}\big),$ (A.53)
with, for the geometrical sector, at the background level,
$\mathcal{H}_{\mathfrak{g}}^{(0)}=-6\sqrt{\mathfrak{p}}g^{2}.$ (A.54)
The first-order term of the geometrical sector reads
$\mathcal{H}_{\mathfrak{g}}^{(1)}=-4\sqrt{\mathfrak{p}}(\tilde{g}+\alpha_{1})\delta K_{b}^{b}-\frac{1}{\sqrt{\mathfrak{p}}}(\tilde{g}^{2}+\alpha_{2})\delta E_{b}^{b}+\frac{2}{\sqrt{\mathfrak{p}}}(1+\alpha_{3})\partial_{a}\partial^{i}\delta E_{i}^{a},$ (A.55)
while the second-order one is
$\mathcal{H}_{\mathfrak{g}}^{(2)}$ $=\sqrt{\mathfrak{p}}(1+\alpha_{4})\delta K_{a}^{b}\delta K^{b}_{a}-\sqrt{\mathfrak{p}}(1+\alpha_{5})(\delta K_{b}^{b})^{2}$
$-\frac{2}{\sqrt{\mathfrak{p}}}(\tilde{g}+\alpha_{6})\delta K^{i}_{a}\delta E^{a}_{i}-\frac{1}{2\mathfrak{p}^{3/2}}(\tilde{g}^{2}+\alpha_{7})\delta E_{b}^{a}\delta E_{a}^{b}$
$+\frac{1}{4\mathfrak{p}^{3/2}}(\tilde{g}^{2}+\alpha_{8})(\delta E_{b}^{b})^{2}+\frac{1}{\mathfrak{p}^{3/2}}(1+\alpha_{9})\mathcal{Z}^{cidj}_{ab}(\partial_{c}\delta E^{a}_{i})(\partial_{d}\delta E^{b}_{j}),$ (A.56)
where
$\mathcal{Z}^{cidj}_{ab}=\frac{1}{4}\epsilon_{k}{}^{ef}\epsilon_{mn}{}^{k}\mathcal{X}^{mjd}_{be}\mathcal{X}^{nic}_{af}-\epsilon_{k}{}^{ie}\mathcal{X}^{kjd}_{be}\delta^{c}_{a}-\epsilon_{k}{}^{ci}\mathcal{X}^{kjd}_{ba}+\frac{1}{2}\delta_{a}^{i}\epsilon_{k}{}^{ce}\mathcal{X}^{kjd}_{be},$ (A.57)
with $\epsilon^{abc}$ the Levi-Civita symbol, $\delta^{a}_{b}$ the Kronnecker delta, while
$\mathcal{X}^{ijb}_{ca}=\epsilon_{c}{}^{ij}\delta_{a}^{b}-\epsilon_{c}{}^{ib}\delta^{j}_{a}+\epsilon^{ijb}\delta_{ca}+\epsilon_{a}{}^{ib}\delta^{j}_{c}.$ (A.58)
For the matter sector of $\mathbb{H}$, the background contribution reads
$\mathcal{H}_{\mathfrak{m}}^{(0)}=\frac{\bm{\pi}^{2}}{2\mathfrak{p}^{3/2}}+\mathfrak{p}^{3/2}V[\bm{\phi}],$ (A.59)
with $\bm{\phi}$ and $\bm{\pi}$ the matter field (with potential $V[\bm{\phi}]$) and its conjugate momentum at the background level. Introducing the perturbation of the scalar field $\delta\phi$ and its conjugate momentum $\delta\pi$, such that $\phi=\bm{\phi}+\delta\phi$ and $\pi=\bm{\pi}+\delta\pi$, the first-order perturbations of the matter sector read
$\mathcal{H}_{\mathfrak{m}}^{(1)}=\frac{\bm{\pi}}{\mathfrak{p}^{3/2}}(1+\beta_{1})\delta\pi-\frac{\bm{\pi}^{2}}{4\mathfrak{p}^{5/2}}(1+\beta_{2})\delta E^{a}_{a}+\mathfrak{p}^{3/2}(\partial_{\phi}V[\bm{\phi}])(1+\beta_{3})\delta\phi+\frac{\sqrt{\mathfrak{p}}}{2}V[\bm{\phi}](1+\beta_{4})\delta E^{a}_{a}.$ (A.60)
The second-order term is
$\mathcal{H}_{\mathfrak{m}}^{(2)}$ $=\frac{1}{2\mathfrak{p}^{3/2}}(1+\beta_{5})(\delta\pi)^{2}-\frac{\bm{\pi}}{2\mathfrak{p}^{5/2}}(1+\beta_{6})\delta\pi\delta E_{a}^{a}$
$+\frac{\mathfrak{p}^{3/2}}{2}(\partial_{\phi}^{2}V[\bm{\phi}])(1+\beta_{7})(\delta\phi)^{2}+\frac{\sqrt{\mathfrak{p}}}{2}(\partial_{\phi}V[\bm{\phi}])(1+\beta_{8})\delta\phi\delta E^{a}_{a}$
$+\frac{\bm{\pi}^{2}}{16\mathfrak{p}^{7/2}}(1+\beta_{9})(\delta E^{a}_{a})^{2}+\frac{\bm{\pi}^{2}}{8\mathfrak{p}^{7/2}}(1+\beta_{10})\delta E^{a}_{b}\delta E^{b}_{a}$
$+\frac{V[\bm{\phi}]}{8\sqrt{\mathfrak{p}}}(1+\beta_{11})(\delta E^{a}_{a})^{2}-\frac{V[\bm{\phi}]}{4\sqrt{\mathfrak{p}}}(1+\beta_{12})\delta E^{a}_{b}\delta E^{b}_{a}$
$+\frac{\sqrt{\mathfrak{p}}}{2}(1+\beta_{13})(\partial_{a}\delta\phi)(\partial^{a}\delta\phi).$ (A.61)
For the diffeomorphism constraint, one gets
$\mathbb{D}\left[N^{a}\right]$ $=\mathbb{D}_{\mathfrak{g}}\left[N^{a}\right]+\mathbb{D}_{\mathfrak{m}}\left[N^{a}\right]$
$=\frac{1}{\kappa}\int d{\bf{x}}\;\delta N^{a}\mathcal{D}^{\mathfrak{g}(1)}_{a}+\frac{1}{\kappa}\int d{\bf{x}}\;\delta N^{a}\mathcal{D}^{\mathfrak{m}(1)}_{a}$, (A.62)
where
$\mathcal{D}^{\mathfrak{g}(1)}_{a}=\mathfrak{p}(1+\xi_{1})\partial_{a}\delta K_{b}^{b}-\mathfrak{p}(1+\xi_{2})\partial_{i}\delta K^{i}_{a}-(h+\xi_{3})\delta^{j}_{a}\partial_{b}\delta E^{b}_{j},$ (A.63)
with $h=h(\mathfrak{c},\mathfrak{p})$ the correction of the curvature $\mathfrak{c}$ in the diffeomorphism constraint. The matter sector reads
$\mathcal{D}^{\mathfrak{m}(1)}_{a}={\bm{\pi}}\partial_{a}\delta\phi.$ (A.64)
Finally, the Gauss constraint is given by
$\mathbb{G}\left[\Lambda^{i}\right]=\frac{1}{\kappa}\int d{\bf{x}}\;\delta\Lambda^{i}\mathcal{G}_{i}^{(1)},$ (A.65)
with
$\mathcal{G}_{i}=(\mathfrak{p}+\chi_{1})\varepsilon_{ij}^{a}\delta K_{a}^{j}+(l+\chi_{2})\varepsilon_{ia}^{j}\delta E^{a}_{j}.$ (A.66)
In section II, the counterterms $\chi_{1}$, $\chi_{2}$, and $\chi_{3}$ are set to zero, while $l=\mathfrak{c}$.
Appendix B List of anomalies without correcting the Gauss constraint
@@
$\mathcal{A}_{1}^{\left\{\mathbb{D},\mathbb{G}\right\}}=(1+\xi_{2})\mathfrak{c}-h-\xi_{3}.$ (B.1)
For the $\left\{\mathbb{H},\mathbb{G}\right\}$ bracket,
$\mathcal{A}_{1}^{\left\{\mathbb{H},\mathbb{G}\right\}}=g^{2}-2\mathfrak{c}(\tilde{g}+\alpha_{6})+(\tilde{g}^{2}+\alpha_{7})+4\mathfrak{p}g\partial_{\mathfrak{p}}g$ (B.2)
$\mathcal{A}_{2}^{\left\{\mathbb{H},\mathbb{G}\right\}}=\mathfrak{c}(1+\alpha_{4})+(\tilde{g}+\alpha_{6})-2g\partial_{\mathfrak{c}}g$ (B.3)
$\mathcal{A}_{3}^{\left\{\mathbb{H},\mathbb{G}\right\}}=-\beta_{10}$ (B.4)
$\mathcal{A}_{4}^{\left\{\mathbb{H},\mathbb{G}\right\}}=\beta_{12}.$ (B.5)
@@
Finally, the $\left\{\mathbb{H},\mathbb{H}\right\}$ bracket,
@@
Appendix C List of anomalies when correcting every constraints
$\mathcal{A}_{1}^{\left\{\mathbb{D},\mathbb{G}\right\}}=$ $\mathfrak{p}(1+\xi_{2})(l+\chi_{2})-(h+\xi_{3})(\mathfrak{p}+\chi_{1}).$ (C.1)
For the $\left\{\mathbb{H},\mathbb{G}\right\}$ bracket, the anomalies read